Ron is 1/2 as old as Sam, who is three times as old as Ted. The sum of their ages is 55. How old is Ron
Answer by richwmiller(17219)

r=s/2
s=3t
r+t+s=55
s/2+t+s=55
3t/2+t+3t=55
multiply by 2
3t+2t+6t=110
11t=110
t=10
s=30
r=15